# Next Tokens Denoising for Speech Synthesis

### v1

@@ Abstract
While diffusion and autoregressive (AR) models have significantly advanced generative modeling, they each present distinct limitations. AR models, which rely on causal attention, cannot exploit future context and suffer from slow generation speeds. Conversely, diffusion models struggle with key-value (KV) caching. To overcome these challenges, we introduce Dragon-FM, a novel text-to-speech (TTS) design that unifies AR and flow-matching. This model processes 48 kHz audio codec tokens in chunks at a compact 12.5 tokens per second rate. This design enables AR modeling across chunks, ensuring global coherence, while parallel flow-matching within chunks facilitates fast iterative denoising. Consequently, the proposed model can utilize KV-cache across chunks and incorporate future context within each chunk. Furthermore, it bridges continuous and discrete feature modeling, demonstrating that continuous AR flow-matching can predict discrete tokens with finite scalar quantizers. This efficient codec and fast chunk-autoregressive architecture also makes the proposed model particularly effective for generating extended content. Experiments [^1] See https://cognitivespeech.github.io/dragon-fm for demos of our work on podcast datasets demonstrate its capability to efficiently generate high-quality zero-shot podcasts.
@@ 1 Introduction
The multi-modal features employed by diffusion and autoregressive (AR) models are commonly categorized as either discrete or continuous. Discrete tokens, whether from byte-pair encoding (BPE) for text or variational autoencoders (VAEs) with quantizers for audio, images, and video, offer advantages in compression and semantic richness, making them well-suited for next-token prediction in language models. A fundamental challenge, however, lies in optimizing training efficiency against compression quality: higher token rates improve reconstruction but significantly complicate long-sequence training for large language models (LLMs), while shorter, more manageable sequences risk compromising reconstruction quality, potentially leading to reduced intelligibility and higher word error rates (WER) in speech. Conversely, continuous features convey richer acoustic and visual details but are inherently more difficult to model. While discrete AR language models perform exceptionally well in text generation due to their manageable BPE-encoded sequence lengths, discrete AR models for speech or images frequently struggle with the extensive token sequences required, often resulting in degraded reconstruction quality. In contrast, diffusion models excel at continuous text-to-image or video generation by effectively leveraging full-context modeling. Yet, both autoregressive and diffusion approaches face a compounding computational burden due to their iterative nature.
@@
Table 1: Comparison between the proposed model and state-of-the-art AR and diffusion-based TTS systems. Here, AR refers to autoregressive modeling, and diffusion indicates diffusion-based models. End-to-end (E2E) denotes joint training of the acoustic model. Low latency indicates small first-byte latency, approaching streaming systems, while low real-time factor (RTF) reflects strong parallel processing capability. We define lower frame rate as smaller than 20 token per second.
In this work, we propose a novel generative modeling and text-to-speech (TTS) framework designed to overcome the limitations of individually optimized autoregressive (AR) and diffusion models. Our framework incorporates a high-quality 48kHz codec that compresses audio into a discrete token sequence at 12.5 tokens per second using a finite scalar quantizer (Mentzer et al., 2023; Parker et al., 2024), ensuring high reconstruction quality and manageable sequence lengths. A key innovation of the proposed model lies in its integration of AR next-token prediction with parallel flow-matching denoising within the same token chunk. This significantly reduces iterative steps while simultaneously preserving AR’s crucial key-value (KV) caching efficiency and leveraging flow-matching’s bidirectional context attention to boost scalability and sampling diversity. Recognizing analogous challenges in text modeling—where long-sequence large language models (LLMs) hinder real-time applications and diffusion models improve performance at an efficiency cost—our acoustic model directly addresses these. It predicts discrete codec tokens via a continuous flow-matching approach, processing 2-second blocks (25 audio tokens) generated simultaneously through flow-matching, with consecutive blocks produced autoregressively, similar to language models. Extensive experiments validate the proposed codec and acoustic model design, powerfully demonstrating the impact of its optimized speech token rates on both latency and performance.
@@ 3 Dragon-FM
The architecture of the model is depicted in Figure 1. Its inference speedup stems from three key factors: (1) the codec sequence is significantly shorter than prior solutions with 48 kHz output; (2) autoregressive iterations are substantially reduced by processing 25 tokens simultaneously, compared to single-token autoregression; and (3) flow-matching denoising steps are minimized through mean flow optimization.
@@ 3.1 Architecture
  • Sequence Length: AR and diffusion transformer (DiT) models rely on Transformer backbones, where shorter sequences reduce latency due to the quadratic complexity of self-attention. It processes a composite sequence comprising the input text, a speaker prompt, and the output speech. To minimize latency, we design a compact 12.5 Hz discrete speech token sequence, which produces high-fidelity 48 kHz audio while maintaining a short sequence length.
@@ 4.1 Datasets
To evaluate our model’s performance and investigate the synergy of autoregressive and diffusion models, we curated a publicly available English podcast dataset comprising 60,000 hours of diverse audio. This dataset encompasses varied topics, speakers, and recording conditions to thoroughly assess the model’s capabilities. An internal automatic speech recognition (ASR) system was used to transcribe the audio, we filter noisy segments, and standardize transcriptions with text frontend rules. Audio data was segmented by speaker, and audio clips with irregular pitch or duration were excluded. Raw audio was not denoised to avoid potential signal loss from denoising models, as the speaker prompt ensures acoustic and timbre consistency between prompt and target audio.
@@ 4.3.1 Diversity evaluation
Podcast generation requires high diversity and natural disfluency, which traditional metrics like pitch variation or duration statistics fail to adequately capture. To assess speech diversity, we use the Fréchet Audio Distance (FAD), adapted from the Fréchet Inception Distance (FID) employed in image generation (Le et al., 2024). FAD quantifies the similarity between the distributions of generated and ground-truth audio, where a smaller distance indicates a smaller gap, reflecting both higher sample quality and greater diversity. As presented in Table 2. Dragon-FM-A3 and Dragon-FM-A1 achieve a higher Fréchet Audio Distance (FAD) score than Dragon-FM-A2, suggesting that additional flow matching steps enhance diversity and similarity, particularly in speech details and fidelity. Dragon-FM-B1 exhibits a lower FAD score than Dragon-FM-A1, likely due to its increased autoregressive loops, which may introduce greater sampling variation. Similarly, Dragon-FM-C1 scores lower than Dragon-FM-A1, despite using more tokens per chunk, possibly because its improved codec reconstruction contributes to a better FAD score.
@@ 4.3.3 Codec reconstruction comparison
Our experimental evaluation, summarized in Table 4, evaluates the performance of our proposed Codec. Unlike traditional Codecs that primarily focus on minimizing bitrate (Bps) or tokens per second, our design objective is to significantly compress the frame rate while preserving high-fidelity audio reconstruction.
We established four baselines for comparison: the original recording and three versions of our internally trained Mel-spectrogram vocoder, operating at frame rates of 25 Hz, 20 Hz, and 16 Hz. The results indicate a clear trend: as the frame rate for the vocoder decreases, there is a corresponding degradation in quality, evidenced by a progressive decline in Speaker Similarity (SIM) and an increase in Word Error Rate (WER). This performance loss becomes particularly pronounced when the frame rate drops below 20 Hz (e.g., the 16 Hz model shows a SIM of 0.904 and WER of 2.80). We attribute this limitation to the short-time stationarity assumption inherent in the Short-Time Fourier Transform (STFT) process, which struggles to maintain quality at lower frame rates.
@@ 5 Conclusion
In this work, we introduced Dragon-FM, a novel text-to-speech (TTS) framework that unifies autoregressive (AR) and flow-matching paradigms to address the fragmented landscape of generative modeling. By integrating a high-quality 48 kHz codec at a compact 12.5 Hz token rate, Dragon-FM achieves efficient acoustic modeling while preserving high-fidelity audio output. Our chunk-wise AR modeling, processing 2-second blocks of 25 tokens, combined with mean flow within each chunk, significantly reduces iteration steps. A key insight from Dragon-FM is the unification of discrete and continuous feature prediction, leveraging continuous vector embeddings of FSQ tokens to harness the intrinsic classification capabilities of continuous denoising models. This approach mitigates the instability of low-frame-rate mel-spectrograms and simplifies training, offering a versatile framework for multimodal generative modeling. By addressing the challenges of sequence length, iteration steps, and feature representation, Dragon-FM not only advances TTS but also provides a scalable solution for text, image, and video generation. Future work will explore extending Dragon-FM’s framework to large-scale multimodal large language models (LLMs), leveraging its unified discrete-continuous approach to enable efficient, high-quality synthesis across diverse data modalities.

### v2

@@ Abstract
While diffusion and autoregressive (AR) models have significantly advanced generative modeling, they each present distinct limitations. AR models, which rely on causal attention, cannot exploit future context and suffer from slow generation speeds. Conversely, diffusion models struggle with key-value (KV) caching. To overcome these challenges, we introduce Dragon-FM, a novel text-to-speech (TTS) design that unifies AR and flow-matching. This model processes 48 kHz audio codec tokens in chunks at a compact rate of 12.5 tokens per second. This design enables AR modeling across chunks, ensuring global coherence, while parallel flow-matching within chunks facilitates fast iterative denoising. Thus, the model leverages KV-cache across chunks and utilizes bidirectional context within each chunk. Furthermore, it bridges continuous and discrete feature modeling, demonstrating that continuous AR flow-matching can predict discrete tokens with finite scalar quantizers. This efficient codec and fast chunk-autoregressive architecture also make the model highly effective for generating long-form content, such as podcasts. Experiments [^1] See https://cognitivespeech.github.io/dragon-fm for demos of our work on podcast datasets demonstrate its capability to efficiently generate high-quality zero-shot podcasts.
@@ 1 Introduction
Features used in diffusion and autoregressive (AR) models for various modalities are commonly categorized as discrete or continuous. Discrete tokens, whether from byte-pair encoding (BPE) for text or variational autoencoders (VAEs) with quantizers for audio, images, and video, offer advantages in compression and semantic richness, making them well-suited for next-token prediction in language models. A fundamental challenge lies in balancing training efficiency with compression quality. Higher token rates improve reconstruction but complicate long-sequence training for large language models (LLMs), while shorter sequences risk compromising reconstruction quality, potentially reducing intelligibility and increasing word error rates (WER) in speech. Conversely, continuous features convey richer acoustic and visual details but are inherently more difficult to model. While discrete AR language models perform exceptionally well in text generation due to their manageable BPE-encoded sequence lengths, discrete AR models for speech or images frequently struggle with the extensive token sequences required, often resulting in degraded reconstruction quality. In contrast, diffusion models excel at continuous text-to-image or video generation by effectively leveraging full-context modeling. Yet, both autoregressive and diffusion approaches face a compounding computational burden due to their iterative nature.
@@
Table 1: Comparison between the proposed model and state-of-the-art AR and diffusion-based TTS systems. Here, AR refers to autoregressive modeling, and diffusion indicates diffusion-based models. End-to-end (E2E) denotes joint training of the acoustic model. Low latency indicates small first-byte latency, approaching streaming systems, while low real-time factor (RTF) reflects strong parallel processing capability. We define a low frame rate as fewer than 15 tokens per second.
In this work, we propose a novel generative modeling and text-to-speech (TTS) framework, Dragon-FM (flow matching), designed to overcome the limitations of individually optimized autoregressive (AR) and diffusion models. Our framework incorporates a high-quality 48kHz codec that compresses audio into a discrete token sequence at 12.5 tokens per second using a finite scalar quantizer (Mentzer et al., 2023; Parker et al., 2024), ensuring high reconstruction quality and manageable sequence lengths. A key innovation of the proposed model lies in its integration of AR next-token prediction with parallel flow-matching denoising within the same token chunk. This significantly reduces iterative steps while simultaneously preserving AR’s crucial key-value (KV) caching efficiency and leveraging flow-matching’s bidirectional context attention to boost scalability and sampling diversity. Recognizing analogous challenges in text modeling, where long-sequence large language models (LLMs) hinder real-time applications and diffusion models improve performance at an efficiency cost, our acoustic model directly addresses these. It predicts discrete codec tokens via a continuous flow-matching approach, processing 2-second blocks (25 audio tokens) generated simultaneously through flow-matching, with consecutive blocks produced autoregressively, similar to language models. We term this approach ’next-token denoising,’ a new paradigm that unifies continuous and discrete features of LLMs and diffusion models, contrasting with traditional discrete token-based next-token prediction. Extensive experiments validate the proposed codec and acoustic model design, powerfully demonstrating the impact of its optimized speech token rates on both latency and performance.
@@ 3 Dragon-FM
The architecture of the model is depicted in Figure 1. Its inference speedup stems from three key factors: (1) the codec sequence is significantly shorter than those of prior solutions for 48 kHz audio output; (2) autoregressive iterations are substantially reduced by processing 25 tokens simultaneously, compared to single-token autoregression; and (3) flow-matching denoising steps are minimized through mean flow optimization.
@@ 3.1 Architecture
  • Sequence Length: AR and diffusion transformer (DiT) models rely on Transformer backbones, where shorter sequences reduce latency due to the quadratic complexity of self-attention. It processes a composite sequence comprising the input text, a speaker prompt, and the output speech. We design a compact discrete speech token sequence at 12.5 tokens per second, which produces high-fidelity 48 kHz audio while maintaining a short sequence length.
@@ 4.1 Datasets
To evaluate our model’s performance and investigate the synergy of autoregressive and diffusion models, we curated a publicly available English podcast dataset comprising 60,000 hours of diverse audio. This dataset encompasses varied topics, speakers, and recording conditions to thoroughly assess the model’s capabilities. An internal automatic speech recognition (ASR) system transcribed the audio, noisy segments were filtered, and transcriptions were standardized using text frontend rules. Audio data was segmented by speaker, and audio clips with irregular pitch or duration were excluded. Raw audio was not denoised to avoid potential signal loss from denoising models, as the speaker prompt ensures acoustic and timbre consistency between prompt and target audio.
@@ 4.3.1 Diversity evaluation
Podcast generation requires high diversity and natural disfluency, which traditional metrics like pitch variation or duration statistics fail to adequately capture. To assess speech diversity, we use the Fréchet Audio Distance (FAD), adapted from the Fréchet Inception Distance (FID) employed in image generation (Le et al., 2024). FAD quantifies the similarity between the distributions of generated and ground-truth audio, where a smaller distance indicates a smaller gap, reflecting both higher sample quality and greater diversity. As presented in Table 2. Dragon-FM-A3 and Dragon-FM-A1 achieve a lower Fréchet Audio Distance (FAD) score than Dragon-FM-A2, suggesting that additional flow-matching steps enhance diversity and similarity, particularly in speech details and fidelity. Dragon-FM-B1 exhibits a lower FAD score than Dragon-FM-A1, likely due to its increased autoregressive loops, which may introduce greater sampling variation. Similarly, Dragon-FM-C1 scores lower than Dragon-FM-A1, despite using more tokens per chunk, possibly because its improved codec reconstruction contributes to a better FAD score.
@@ 4.3.3 Codec reconstruction comparison
Our experimental evaluation, summarized in Table 4, evaluates the performance of our proposed Codec. Unlike traditional codecs that primarily focus on minimizing bitrate (bits/s) or token rate, our design objective is to significantly compress the frame rate while preserving high-fidelity audio reconstruction.
We established four baselines for comparison: the original recording and three versions of our internally trained Mel-spectrogram vocoder, operating at frame rates of 25 Hz, 20 Hz, and 16 Hz. The results indicate a clear trend: as the frame rate for the vocoder decreases, there is a corresponding degradation in quality, evidenced by a progressive decline in Speaker Similarity (SIM) and an increase in Word Error Rate (WER). This performance loss becomes particularly pronounced when the frame rate drops below 20 Hz (e.g., the 16 Hz model shows a SIM of 0.904 and WER of 3.01). We attribute this limitation to the short-time stationarity assumption inherent in the Short-Time Fourier Transform (STFT) process, which struggles to maintain quality at lower frame rates.
@@ 5 Conclusion
In this work, we introduced Dragon-FM, a novel text-to-speech (TTS) framework that unifies autoregressive (AR) and flow-matching paradigms to address the fragmented landscape of generative modeling. By integrating a high-quality 48 kHz codec at a compact 12.5 Hz token rate, Dragon-FM achieves efficient acoustic modeling while preserving high-fidelity audio output. Our chunk-wise AR modeling, processing 2-second blocks of 25 tokens, combined with mean flow within each chunk, significantly reduces iteration steps. A key insight from Dragon-FM is the unification of discrete and continuous feature prediction, leveraging continuous vector embeddings of FSQ tokens to harness the intrinsic classification capabilities of continuous denoising models. This approach mitigates the instability of low-frame-rate mel-spectrograms and simplifies training, offering a versatile framework for multimodal generative modeling. By addressing the challenges of sequence length, iteration steps, and feature representation, Dragon-FM advances TTS and offers a framework potentially adaptable to other generative tasks. Future work will explore extending Dragon-FM’s framework to large-scale multimodal large language models (LLMs), leveraging its unified discrete-continuous approach to enable efficient, high-quality synthesis across diverse data modalities.